# Exploring Inevitable Waypoints for Unsolvability Explanation in Hybrid Planning Problems

### v1

@@ 4.3. Explanation Generation by Reachability Analysis
We use a bounded reachability analysis tool Bach (DBLP:conf/fmcad/BuLWL08, ) for reachability analysis of the waypoints. Bach can analyse linear hybrid automata (LHA) (DBLP:conf/lics/Henzinger96, ; DBLP:journals/entcs/LiAB07, ) and reports a reachability problem instance as satisfiable when a run exists from an initial state to a goal state in the corresponding hybrid automaton for the planning problem of length bounded by a given depth, deciding the solvability of the corresponding planning problem. Otherwise, Bach reports the instance as unsatisfiable, when no such run exists. Algorithm 1 takes a chain of sub-problems $Y^{<}$ as input. It returns the first unreachable sub-problem in the chain as an explanation of unsolvability.
@@
       $\mathcal{R}\leftarrow$ reachabilityProblem(HA, $Init$, $Goal_{l_{k}}$, $Depth$), where $Goal_{l_{k}}$ is the goal of $\Pi_{l_{k}}$
@@ 4.4. Complexity Analysis
In a HA with $n$ locations and given a source and destination location, the worst-case complexity of computing all paths in $\mathcal{G}_{HA}$ of length at most $d$ from the given source to the destination is $O\big{(}n^{d}\big{)}$. The complexity of computing $LCS$ of strings corresponding to the paths (say $m$ many) of maximum length $d$ using standard dynamic programming is $O\big{(}d^{m}\big{)}$. It is known to be an NP-Hard problem (DBLP:journals/asc/DjukanovicRB20, ; DBLP:journals/jacm/Maier78, ). Although finding the inevitable way-points with the proposed algorithm turns out to be inefficient asymptotically, for problem instances of small size (HA with a few locations and for a small depth d), we show empirically that our algorithm can generate inevitable way-points and the explanation artifact efficiently.
@@
The model checker Bach performs a path-oriented reachability analysis of a linear hybrid automaton. An initial to goal location path is encoded into a set of linear constraints, and consequently solved using linear programming problem solver. Details of the path encoding can be found in (DBLP:conf/fmcad/BuLWL08, ; DBLP:conf/memocode/SarwarRB23, ). Though practical LP solvers use variants of the Simplex algorithm, which runs efficiently on most practical problems, it is not a polynomial time algorithm in general. In theory, linear programming has been shown to be in class P (khachiyan1979polynomial, ; KHACHIYAN198053, ). Our algorithm calls Bach for each sub-problem in the computed chain. As the length of the computed chain, the LCS, is bounded by the number of locations of HA (proposition 4.10), the complexity is $|\Pi^{*}|$ times the complexity of LP solving.
@@ 5.1. Experimental Setup
: A brief description of the planning domains is given as follows: Planetary rover domain is presented in Section 2 and a pictorial overview is shown in Figure 1. The planning task for the rover is to reach the base station from its initial location after collecting soil and rock samples from the designated sites. City route-network domain presents a route-network of a city where important places are given as junctions in the network. The planning problem for a battery-powered car is to navigate through the city’s route network to reach its destination. Warehouse automation domain (DBLP:conf/memocode/SarwarRB23, ) represents a scenario where a robot operates to manage the inventories of a warehouse. The floor map of the warehouse is given as grid cells. Few cells in the warehouse are blocked, whereas, on a few cells, the robot depletes more energy due to the condition of the surface such as a oil spillage or being bumpy. The planning problem is for the robot to carry a consignment from its initial location to a goal location. We have crafted warehouse scenarios of varying grid dimensions for evaluating our algorithm. Water-level monitor (DBLP:conf/fmcad/BuLWL08, ) represents a system that controls the water level in a reservoir. The system goes into an unsafe state if the water level in the reservoir meets underflow or overflow conditions. The planning task is to drive the system to an unsafe state from a given initial state. NAV (ARCH-COMP24:ARCH_COMP24_Category_Report, ) models the motion of a point robot in a $2$-dimensional plane, partitioned into $3^{2}$ rectangular regions, and each such region is associated with a vector field described by the flow equations. The planning problem is to find a trajectory from an initial state to a goal state. NRS (F-Wang-NRS, ; ARCH-COMP24:ARCH_COMP24_Category_Report, ) represents a nuclear reactor system consisting of 2 rods that absorb neutrons from heavy water when inserted, and a controller that schedules the insertion of the rods into the heavy water. The system is considered safe if there is exactly one rod absorbing neutrons in the heavy water at any instant of time. The planning problem is to find an unsafe execution of the system from a given initial state. A detailed description of the city route and warehouse automation domains is given in the Appendix.
We constructed the Planetary rover and the City route-network domains for evaluating our algorithm. The Warehouse automation domain is taken from (DBLP:conf/memocode/SarwarRB23, ). The rest of the domains are from verification problem instances in linear hybrid systems. For instance, the Water-level monitor domain is a benchmark taken from (DBLP:conf/fmcad/BuLWL08, ), whereas NAV and NRS are benchmarks taken from Arch-comp 24 pcdb category  (ARCH-COMP24:ARCH_COMP24_Category_Report, ). In these domains, we pose the safety property verification problem as planning problem instances. The planning problems taken for evaluation are all known to be unsolvable.
@@ 5.3. Analysis of results
In Figure 6, we illustrate our method in the example scenarios of Planetry rover and Warehouse automation domains. In the motivating example problem instance in Planetary rover domain, the algorithm identified 7 sub-problems $\{W1,W2,\ldots,W7\}$ in Y< with the total order $W1<W2<\ldots<W7$, each representing an inevitable waypoint. Our explanation algorithm detects five waypoints ($W1$-$W5$) as reachable depicted by Green ticks and reports $W6$ as the first unreachable waypoint shown by a Red cross in Figure 6(a). The unreachability of $W6$ can lead the human expert to deduce that the rover’s initial battery charge is insufficient to drive it past the ascending regions in $W4$ and $W5$. Figure 6(b) shows the identified waypoints and an explanation on a 6$\times$6 warehouse domain for a planning problem where a robot needs to carry a consignment from its initial location to the goal location. In the domain, Blue and Red cells are the initial and goal locations, respectively. Yellow cells have surfaces with oil-spillage and therefore robot has a greater rate of battery depletion in these cells. Grey cells are blocked. The Green cell is the only charging station. The explanation algorithm identified 6 sub-problems $\{W1,W2,\ldots,W6\}$ for a planning problem depth bound of 12, and 4 sub-problems $W1$ to $W4$ for a depth bound of 17, respectively. The waypoints $W5$ and $W6$ are inevitable only when the depth bound is 12 since a path longer than 12 in length may not mandatorily visit these waypoints. Note that these vertices ($W5$ and $W6$) are not articulation points whereas the other waypoints ($W1-W4$) are articulation points of the warehouse grid graph. In both the problem instances, our algorithm reports that the robot can not reach the waypoint $W4$ under the dynamics which is the an explanation of unsolvability. A control engineer can deduce that the initial battery charge and the charge capacity of the robot is not sufficient to reach the waypoint directly or via the recharging station. Therefore, a higher charge capacity or a better placement of the charging station close to the waypoint $W4$ may be a work-around to make the task solvable.
@@ 6. Related works
Some notable works addressing the unsolvability of planning problems, mostly looked at verifying the unsolvability by generating certificates (DBLP:conf/aips/ErikssonH20, ), (DBLP:conf/aips/ErikssonRH17, ) or proofs (DBLP:conf/aips/ErikssonRH18, ) rather than explaining the causalities of unsolvability of the planning problem. Such certificates or proofs of unsolvability are not enough to increase the human understandability of why the problem was unsolvable. Most of these works focus on planning problems in discrete domains. Verifying the unsolvability of planning problems in hybrid systems comes with an additional challenge since these planning problems are undecidable in general (ALUR19953, ). In (DBLP:journals/tecs/SarwarRB23, ), authors provide an approach to addressing the unsolvability of a planning problem in hybrid domains by $\delta$-approximate bounded reachability analysis (DBLP:journals/corr/GaoKCC14, ). However, this work also verifies unsolvability rather than explaining it. Few notable works that are directed towards explaining the unsolvability of a planning problem are, similarly, limited to classical planning problems. Authors in (DBLP:conf/aips/GobelbeckerKEBN10, ) argue that excuses can be produced by counterfactual alterations to the original planning task such that the new planning task turns out to be solvable, and provides excuses for why a plan cannot be found. In (DBLP:conf/aaai/EiflerC0MS20, ), authors derive properties of a plan which could serve as explanations in case of unsolvability. However, generating excuses, or deriving plan properties in terms of propositional formulas may not be enough to understand why a problem was unsolvable for complex domains like planning problems of hybrid systems which encode mixed discrete and continuous dynamics. In (DBLP:journals/jair/VasileiouYSKCM22, ), an approach based on knowledge representation and reasoning has been applied to these domains. It provides explanations by finding a subset of the agent’s knowledge base with which to reconcile the human knowledge base for explanations. However, it does not address unsolvability problems, rather, explains why a plan is feasible in a model. A path-oriented reconciliation process between the agent and human models of hybrid systems is provided in (DBLP:conf/memocode/SarwarRB23, ). It performs the reachability analysis along a path and uses the concept of irreducible infeasible sets (IIS) to generate explanations for unsolvability.
@@
In this work, we propose to decompose an unsolvable planning problem into sub-problems motivated by the well-known insight that humans tend to break down sequential planning problems in terms of the sub-problems they need to achieve (LT-1956, ; DBLP:journals/ai/VanLehn86, ). This has been a popular approach in many domains such as robotics (24200, ) and AI (DBLP:journals/ai/SuttonPS99, ) apart from planning (DBLP:journals/jair/HoffmannPS04, ; DBLP:conf/ecai/LipovetzkyG12, ; DBLP:conf/aaai/RichterHW08, ). (DBLP:journals/jair/HoffmannPS04, ; DBLP:conf/ecai/LipovetzkyG12, ) find sub-problems for a solvable planning problem of the discrete domains in terms of ordered landmarks. Landmarks are facts given as propositional formulas that must be true at some point in every valid solution plan. In (DBLP:conf/ijcai/SreedharanSSK19, ), authors use hierarchical model abstractions to relax a planning problem until a solution can be found and looks for landmarks of this relaxed problem. They use these landmarks to identify the unachievable sub-problem for the planning problem. These works are in discrete domains. In contrast, our framework decomposes an unsolvable planning problem of hybrid domains into several smaller sub-problems by reducing it to an instance of longest common subsequence problem and consequently generating explanations using reachability analysis.

### v2

@@ 4.3. Explanation Generation by Reachability Analysis
We use a bounded reachability analysis tool Bach (DBLP:conf/fmcad/BuLWL08) for reachability analysis of the waypoints. Bach can analyse linear hybrid automata (LHA) (DBLP:conf/lics/Henzinger96; DBLP:journals/entcs/LiAB07) and reports a reachability problem instance as satisfiable when a run exists from an initial state to a goal state in the corresponding hybrid automaton for the planning problem of length bounded by a given depth, deciding the solvability of the corresponding planning problem. Otherwise, Bach reports the instance as unsatisfiable, when no such run exists. Algorithm 1 takes a chain of sub-problems $Y^{<}$ as input. It returns the first unreachable sub-problem in the chain as an explanation of unsolvability.
@@
  $\mathcal{R}\leftarrow$ reachabilityProblem(HA, InitInit, $Goal_{l_{k}}$, DepthDepth), where $Goal_{l_{k}}$ is the goal of $\Pi_{l_{k}}$
@@ 4.4. Complexity Analysis
In a HA with $n$ locations and given a source and destination location, the worst-case complexity of computing all paths in $\mathcal{G}_{HA}$ of length at most $d$ from the given source to the destination is $O\big(n^{d}\big)$. The complexity of computing $LCS$ of strings corresponding to the paths (say $m$ many) of maximum length $d$ using standard dynamic programming is $O\big(d^{m}\big)$. It is known to be an NP-Hard problem (DBLP:journals/asc/DjukanovicRB20; DBLP:journals/jacm/Maier78). Although finding the inevitable way-points with the proposed algorithm turns out to be inefficient asymptotically, for problem instances of small size (HA with a few locations and for a small depth d), we show empirically that our algorithm can generate inevitable way-points and the explanation artifact efficiently.
@@
The model checker Bach performs a path-oriented reachability analysis of a linear hybrid automaton. An initial to goal location path is encoded into a set of linear constraints, and consequently solved using linear programming problem solver. Details of the path encoding can be found in (DBLP:conf/fmcad/BuLWL08; DBLP:conf/memocode/SarwarRB23). Though practical LP solvers use variants of the Simplex algorithm, which runs efficiently on most practical problems, it is not a polynomial time algorithm in general. In theory, linear programming has been shown to be in class P (khachiyan1979polynomial; KHACHIYAN198053). Our algorithm calls Bach for each sub-problem in the computed chain. As the length of the computed chain, the LCS, is bounded by the number of locations of HA (proposition 4.10), the complexity is $|\Pi^{*}|$ times the complexity of LP solving.
@@ 5.1. Experimental Setup
: A brief description of the planning domains is given as follows: Planetary rover domain is presented in Section 2 and a pictorial overview is shown in Figure 1. The planning task for the rover is to reach the base station from its initial location after collecting soil and rock samples from the designated sites. City route-network domain presents a route-network of a city where important places are given as junctions in the network. The planning problem for a battery-powered car is to navigate through the city’s route network to reach its destination. Warehouse automation domain (DBLP:conf/memocode/SarwarRB23) represents a scenario where a robot operates to manage the inventories of a warehouse. The floor map of the warehouse is given as grid cells. Few cells in the warehouse are blocked, whereas, on a few cells, the robot depletes more energy due to the condition of the surface such as a oil spillage or being bumpy. The planning problem is for the robot to carry a consignment from its initial location to a goal location. We have crafted warehouse scenarios of varying grid dimensions for evaluating our algorithm. Water-level monitor (DBLP:conf/fmcad/BuLWL08) represents a system that controls the water level in a reservoir. The system goes into an unsafe state if the water level in the reservoir meets underflow or overflow conditions. The planning task is to drive the system to an unsafe state from a given initial state. NAV (ARCH-COMP24:ARCH_COMP24_Category_Report) models the motion of a point robot in a $2$-dimensional plane, partitioned into $3^{2}$ rectangular regions, and each such region is associated with a vector field described by the flow equations. The planning problem is to find a trajectory from an initial state to a goal state. NRS (F-Wang-NRS; ARCH-COMP24:ARCH_COMP24_Category_Report) represents a nuclear reactor system consisting of 2 rods that absorb neutrons from heavy water when inserted, and a controller that schedules the insertion of the rods into the heavy water. The system is considered safe if there is exactly one rod absorbing neutrons in the heavy water at any instant of time. The planning problem is to find an unsafe execution of the system from a given initial state. A detailed description of the city route and warehouse automation domains is given in the Appendix.
We constructed the Planetary rover and the City route-network domains for evaluating our algorithm. The Warehouse automation domain is taken from (DBLP:conf/memocode/SarwarRB23). The rest of the domains are from verification problem instances in linear hybrid systems. For instance, the Water-level monitor domain is a benchmark taken from (DBLP:conf/fmcad/BuLWL08), whereas NAV and NRS are benchmarks taken from Arch-comp 24 pcdb category  (ARCH-COMP24:ARCH_COMP24_Category_Report). In these domains, we pose the safety property verification problem as planning problem instances. The planning problems taken for evaluation are all known to be unsolvable.
@@ 5.3. Analysis of results
In Figure 6, we illustrate our method in the example scenarios of Planetry rover and Warehouse automation domains. In the motivating example problem instance in Planetary rover domain, the algorithm identified 7 sub-problems $\{W1,W2,\ldots,W7\}$ in Y< with the total order $W1<W2<\ldots<W7$, each representing an inevitable waypoint. Our explanation algorithm detects five waypoints ($W1$-$W5$) as reachable depicted by Green ticks and reports $W6$ as the first unreachable waypoint shown by a Red cross in Figure 6(a). The unreachability of $W6$ can lead the human expert to deduce that the rover’s initial battery charge is insufficient to drive it past the ascending regions in $W4$ and $W5$. Figure 6(b) shows the identified waypoints and an explanation on a $\times6$ warehouse domain for a planning problem where a robot needs to carry a consignment from its initial location to the goal location. In the domain, Blue and Red cells are the initial and goal locations, respectively. Yellow cells have surfaces with oil-spillage and therefore robot has a greater rate of battery depletion in these cells. Grey cells are blocked. The Green cell is the only charging station. The explanation algorithm identified 6 sub-problems $\{W1,W2,\ldots,W6\}$ for a planning problem depth bound of 12, and 4 sub-problems $W1$ to $W4$ for a depth bound of 17, respectively. The waypoints $W5$ and $W6$ are inevitable only when the depth bound is 12 since a path longer than 12 in length may not mandatorily visit these waypoints. Note that these vertices ($W5$ and $W6$) are not articulation points whereas the other waypoints (W1W4W1-W4) are articulation points of the warehouse grid graph. In both the problem instances, our algorithm reports that the robot can not reach the waypoint $W4$ under the dynamics which is the an explanation of unsolvability. A control engineer can deduce that the initial battery charge and the charge capacity of the robot is not sufficient to reach the waypoint directly or via the recharging station. Therefore, a higher charge capacity or a better placement of the charging station close to the waypoint $W4$ may be a work-around to make the task solvable.
@@ 6. Related works
Some notable works addressing the unsolvability of planning problems, mostly looked at verifying the unsolvability by generating certificates (DBLP:conf/aips/ErikssonH20), (DBLP:conf/aips/ErikssonRH17) or proofs (DBLP:conf/aips/ErikssonRH18) rather than explaining the causalities of unsolvability of the planning problem. Such certificates or proofs of unsolvability are not enough to increase the human understandability of why the problem was unsolvable. Most of these works focus on planning problems in discrete domains. Verifying the unsolvability of planning problems in hybrid systems comes with an additional challenge since these planning problems are undecidable in general (ALUR19953). In (DBLP:journals/tecs/SarwarRB23), authors provide an approach to addressing the unsolvability of a planning problem in hybrid domains by $\delta$-approximate bounded reachability analysis (DBLP:journals/corr/GaoKCC14). However, this work also verifies unsolvability rather than explaining it. Few notable works that are directed towards explaining the unsolvability of a planning problem are, similarly, limited to classical planning problems. Authors in (DBLP:conf/aips/GobelbeckerKEBN10) argue that excuses can be produced by counterfactual alterations to the original planning task such that the new planning task turns out to be solvable, and provides excuses for why a plan cannot be found. In (DBLP:conf/aaai/EiflerC0MS20), authors derive properties of a plan which could serve as explanations in case of unsolvability. However, generating excuses, or deriving plan properties in terms of propositional formulas may not be enough to understand why a problem was unsolvable for complex domains like planning problems of hybrid systems which encode mixed discrete and continuous dynamics. In (DBLP:journals/jair/VasileiouYSKCM22), an approach based on knowledge representation and reasoning has been applied to these domains. It provides explanations by finding a subset of the agent’s knowledge base with which to reconcile the human knowledge base for explanations. However, it does not address unsolvability problems, rather, explains why a plan is feasible in a model. A path-oriented reconciliation process between the agent and human models of hybrid systems is provided in (DBLP:conf/memocode/SarwarRB23). It performs the reachability analysis along a path and uses the concept of irreducible infeasible sets (IIS) to generate explanations for unsolvability.
@@
In this work, we propose to decompose an unsolvable planning problem into sub-problems motivated by the well-known insight that humans tend to break down sequential planning problems in terms of the sub-problems they need to achieve (LT-1956; DBLP:journals/ai/VanLehn86). This has been a popular approach in many domains such as robotics (24200) and AI (DBLP:journals/ai/SuttonPS99) apart from planning (DBLP:journals/jair/HoffmannPS04; DBLP:conf/ecai/LipovetzkyG12; DBLP:conf/aaai/RichterHW08). (DBLP:journals/jair/HoffmannPS04; DBLP:conf/ecai/LipovetzkyG12) find sub-problems for a solvable planning problem of the discrete domains in terms of ordered landmarks. Landmarks are facts given as propositional formulas that must be true at some point in every valid solution plan. In (DBLP:conf/ijcai/SreedharanSSK19), authors use hierarchical model abstractions to relax a planning problem until a solution can be found and looks for landmarks of this relaxed problem. They use these landmarks to identify the unachievable sub-problem for the planning problem. These works are in discrete domains. In contrast, our framework decomposes an unsolvable planning problem of hybrid domains into several smaller sub-problems by reducing it to an instance of longest common subsequence problem and consequently generating explanations using reachability analysis.
@@ 7. Conclusion
8. Appendix
City-network domain: the context for this domain is a car that wants to reach a destination through a route network of a city. Figure 7 shows the route network of the city. It has 10 important junctures. Blue-colored and green-colored routes connect these junctures. They respectively represent both-way and one-way traffic in the city. The direction of the traffic in green routes is shown with a directed arrow. The car is initially at juncture A. The juncture A is shown as the green-colored node in the figure. The car has a battery that depletes energy at a constant rate represented by a variable $b$. Initially, it has 20 units of battery charges available. Similarly, the juncture-to-juncture movement delay for the car is represented by a variable $d$ where each route has a different delay. From a juncture, the car can only move to the adjacent junctures following the route between them. The junctures H (red-colored), I (yellow-colored), and J (blue-colored) are the destinations of three different planning problems of the domain. The orange-colored nodes in the figure are the waypoints that appear in every source-to-destination path for a planning problem of A to H. The routes represent the discrete dynamics of the domain that captures the connectivity of the junctures of the city. The continuous dynamics of the domain involve energy depletion and the juncture-to-juncture movement delay of the car due to different traffic patterns.
Figure 7. The city route network is depicted. Each node represents the important junctures of the city. The blue-colored routes represent both-way transportation between junctures. The green-colored routes represent one-way traffic. Initially, the car is at juncture A (shown as the green-colored node). The junctures I (yellow-colored), H (red-colored), and J (blue-colored) are the destinations of three different planning problems of the domain. The orange-colored nodes in the figure are the waypoints that appear in every source-to-destination path for a planning problem of A to H.
Warehouse automation domain: We present the context of warehouse automation [DBLP:conf/memocode/SarwarRB23] where a robot operates to manage the inventories of the warehouse. The warehouse is divided into cells. The discrete dynamics here capture the connectivity of the cells along with the presence of objects in certain cells which are interpreted as obstacles through which the robot cannot move. The movement of the robot is restricted to one of its adjacent cells and movement to diagonal cells is prohibited. The continuous dynamics capture the battery charge depletion rate of the robot within a cell. Within each cell, the robot follows the dynamics particular to that cell. When the robot makes a transition from one cell to another, it starts to follow the dynamics of the new cell instantaneously. The robot is assigned the task of carrying a consignment to a designated cell while the number of cell visits is restricted to $\leq$ $D$ cells. The robot starts from the yellow-colored cell where the planning problem requires it to transport the black box to the goal cell (red-colored cell). The robot depletes its charge according to the cell dynamics while on the move. There is a charging station shown as a green-colored cell. The robot may visit this cell to recharge its battery. The grey-colored cells are blocked with obstacles. The robot is equipped with a rechargeable battery. The initial battery charge is 10 units. Each cell has a charge depletion rate of 2 units (modeling the continuous dynamics). In Figure 8, we have shown a representation of the warehouse automation domain. Now, consider the planning problem where the robot needs to carry the consignment to the goal from its initial location. Every feasible path for the robot must go through the cells marked with hatched lines (orange-colored) as shown in the figure. We consider these cells as landmarks for the planning problem. A landmark here means a cell that a robot must visit on its way to the goal.
Figure 8. Warehouse automation domain.